# Inferring Dynamic Networks from Marginals with Iterative Proportional Fitting

### v1

@@ B.3 The “joint” network inference problem
for all $i,j$ where $\bar{X}_{ij}>0$; otherwise, $X_{ij}=0$ or $\sum_{t}e^{-v_{it}+u_{jt}}=0$. In other words, the aggregated time-dependent intensities for each pair $(i,j)$ does not depend on $i,j$. We may justify (17) in several ways. For example, if we assume $v_{it}$ and $u_{jt}$ are drawn i.i.d. from two distributions with the same mean, then $\frac{1}{T}\sum_{t}e^{-v_{it}+u_{jt}}\rightarrow 1$ by the continuous mapping theorem. Alternatively, if $X_{ij}$ is an accurate description of the the long-run traffic propensity between $i,j$, then the intensity of the aggregate traffic Poisson variables $\sum_{t}Y_{ijt}$ should largely be captured by $X_{ij}$, i.e., independent of $\sum_{t}e^{-v_{it}+u_{jt}}$.
B.4 When IPF recovers the true network exactly
@@
Second, we prove that, if $X^{(t)}=A\bar{X}B$, then IPF will return $X^{(t)}$. First, if there is a finite solution to the matrix balancing problem, IPF will converge to a solution (Pukelsheim, 2014). We know that $A\bar{X}B$ is a solution, so IPF will return $D^{0}\bar{X}D^{1}$, where $(D^{0}\bar{X}D^{1})\cdot\mathbf{1}_{n}=p^{(t)}$ and $(D^{1}\bar{X}^{T}D^{0})\cdot\mathbf{1}_{m}=q^{(t)}$. Furthermore, biproportional scalings are unique with respect to marginals, meaning meaning if two biproportional scalings $M^{1}$ and $M^{2}$ of $M^{0}$ have the same marginals, then $M^{1}=M^{2}$ (Pukelsheim, 2014). Thus, IPF will return $X^{(t)}$, since IPF will converge to $D^{0}\bar{X}D^{1}$, with marginals $p^{(t)}$ and $q^{(t)}$, and $X^{(t)}$ is the unique biproportional scaling of $\bar{X}$ that matches the marginals. ∎
@@
B.5 Poisson links uniquely recover IPF within a class of generalized linear models
@@ C.1.1 Concentration inequalities with Orlicz norms
[…]
@@ C.2 Well-posedness of maximum likelihood estimation of biproportional Poisson
Recall that the negative log-likelihood of the biproportional Poisson model is equivalent to
@@
Note that each $Y_{ij}$ is an independent Poisson random variable with mean and variance parameter $e^{u_{i}}X_{ij}e^{-v_{j}}$. Now for each non-empty $I\subseteq[m]$ and $J\subseteq[n]$ with $X_{IJ^{C}}=0$, i.e., $X_{ij}\equiv 0$ for any $(i,j)$ satisfying $i\in I^{C},j\in J$, let $E_{IJ}$ be the event that $\sum_{j\in J}q_{j}=\sum_{i\in I}p_{i}$. The event $E_{IJ}$ corresponds to the case where the realized network of Poisson variables being disconnected, hence the MLE being unbounded. Our goal is to bound the union of all events $E_{IJ}$ for $X_{IJ^{C}}=0$. We have
@@
where in the last step we used Bennett’s inequality (Bennett, 1962).
@@
where we have used the fact that $X_{ij}\equiv 0$ for any $(i,j)$ satisfying $i\in I^{C},j\in J$. Now using once again that $\mathbf{1}_{m+n}$ spans the null space of $\mathcal{L}(u,v)$, we have
$\sum_{i\in I,j\in J^{C}}e^{u_{i}}X_{ij}e^{-v_{j}}$ $=\begin{bmatrix}e_{I}\\ e_{J}\end{bmatrix}^{T}\mathcal{L}(u^{\ast},v^{\ast})\begin{bmatrix}e_{I}\\ e_{J}\end{bmatrix}$
@@
Now applying a union bound on all $I,J$ with $X_{IJ^{C}}=0$, we have
(MLE does not $\mathbb{P}(\text{MLE does not exist})$ $\leq\sum_{I\subsetneq[m],J\subsetneq[n],X_{IJ^{C}}=0}\exp(-\frac{1}{2}\lambda_{-2}(\mathcal{L}(u^{\ast},v^{\ast}))\frac{(|I|+|J|)(m+n-|I|-|J|)}{m+n})$
@@
(MLE does not $\mathbb{P}(\text{MLE does not exist})$ $\leq 2\left[\left(1+\exp(-\frac{1}{2}\log(m+n))\right)^{m+n}-1\right]$
$=2\left[\left(1+(m+n)^{-1/2}\right)^{m+n}-1\right]$
@@ E.1.2 Evaluating IPF over varying sparsity in $\bar{X}$
  • The bound on the MLE’s expected estimation error, $\mathbb{E}[||(\hat{u}-u,\hat{v}-v)||^{2}_{2}]$, from Theorem 4.1, without the constant:
@@
Figure E.3: Comparing sparsity rate in $\bar{X}$ to number of IPF iterations (left), bound on MLE’s expected estimation error (middle), and observed $\ell_{2}$ error of IPF estimates (right). Lines represent mean and shaded region represents 95% CIs, over 1000 trials per sparsity rate.
In Figure E.3, we visualize our results, showing the mean and 95% CIs (from 2.5th to 97.5th percentiles over random trials) per metric over sparsity rates. We find that IPF is negatively affected by sparsity in multiple ways: the number of iterations until convergence increases, and both the expected bound and observed estimation error on the network parameters worsens. However, the rate at which these metrics change differ, with the bound growing most quickly at high levels of sparsity. To better understand why, below we analyze the connection between the bound on the MLE’s expected error, which we derived in Theorem 4.1, and the observed $\ell_{2}$ error that we evaluate in these experiments.
@@
and the right hand side qualitatively matches the shape observed in Figure E.3 (right).
Sketch of concentration argument. We only include a sketch of this argument since it is standard in the random graph/matrix literature and not part of any of the main results of this work. By standard concentration estimates (i.e. Bernstein’s inequality), the number of entries of $X$ that are deleted from each row will be $rn\pm O(\sqrt{rn\log(n)})$. Equivalently, the number of entries that remain in each row will be $(1-r)n\pm O(\sqrt{rn\log(n)})$. Next we can compute that that
@@ E.2.2 Convergence experiments on mobility data
Replace all zeros
ConvIPF, min # edges, largest $p_{i}$
ConvIPF, min # edges, smallest $p_{i}$
ConvIPF, min change in $\lambda_{1}$
@@
# new edges
$10012193$
$\mathbf{2}$
$16$
$5$
Change in $\lambda_{1}$
$30.04$
$1.85\cdot 10^{-8}$
$2.54\cdot 10^{-7}$
$\mathbf{5.60\cdot 10^{-9}}$
# ConvIPF iters.
$\mathbf{2}$
$16$
$5$
@@
# new edges
$10012193$
$\mathbf{2}$
$3$
$3$
Change in $\lambda_{1}$
$30.04$
$2.14\cdot 10^{-7}$
$1.02\cdot 10^{-8}$
$\mathbf{9.98\cdot 10^{-9}}$
# ConvIPF iters.
$\mathbf{1}$
$2$
$2$
@@
# new edges
$10012193$
$\mathbf{1}$
$9$
$5$
Change in $\lambda_{1}$
$30.04$
$1.63\cdot 10^{-8}$
$8.52\cdot 10^{-8}$
$\mathbf{5.72\cdot 10^{-9}}$
# ConvIPF iters.
$\mathbf{1}$
$9$
$5$
@@ E.3 Experiments with bikeshare data
The distribution of trips over stations is highly skewed (Figure E.11). To aggregate the individual trips into hourly network, we need to assign each trip to an hour. In 79% of trips, the trip starts and ends in the same hour, so the assignment is unambiguous. For the remaining 21% of trips, we assign the trip to the hour of its midway point. So, in our hourly network $X^{(t)}$, each $X^{(t)}_{ij}$ represents the total number of trips from station $i$ to station $j$ with its midpoint falling in hour $t$.
Figure E.12: Distribution of bike trips over hours, from September 1 12AM to September 30 11PM, 2023.
In Figure E.12, we plot the distribution of trips over hours. We can see clear temporal patterns, e.g., fewer trips at night, two peaks on weekdays for getting to and getting from work. As usual, $p^{(t)}$ and $q^{(t)}$ are the row sums and column sums, respectively, of $X^{(t)}$. For the time-aggregated network $\bar{X}$, we explore different temporal aggregations (month, week, and day), so that we can experiment with IPF performance as the input network becomes increasingly time-aggregated.
@@ E.3.2 Convergence behavior
We observe a tradeoff between accuracy and convergence time. Using the day-aggregated network for IPF consistently outperforms the week-aggregated or month-aggregated network at recovering the true hourly network (Table E.3), but running IPF on the day-aggregated network takes substantially more iterations than using the week-aggregated network or month-aggregated network. Over the 24 hours we tested, the mean and median number of iterations when using the day-aggregated network were 7643.29 and 10000, respectively, while they were 920.79 and 419 for the week-aggregated network and 537.83 and 372 for the month-aggregated network. [^8] These are actually underestimates, since we used a maximum of 10,000 iterations. Using the month-aggregated network never reached this maximum, week-aggregated reached it in one hour out of 24, and day-aggregated reached it in most hours. This is not surprising, since the day-aggregated network is far sparser (2.04% nonzero) than the week-aggregated network (7.05% nonzero) or month-aggregated network (13.72% nonzero), and we know that IPF iterations increase with sparsity, which we also saw with synthetic data (Figure E.3, left).

### v2

@@ B.4 The “joint” network inference problem
for all $i,j$ where $\bar{X}_{ij}>0$; otherwise, $X_{ij}=0$ or $\sum_{t}e^{-v_{it}+u_{jt}}=0$. In other words, the aggregated time-dependent intensities for each pair $(i,j)$ does not depend on $i,j$. We may justify (17) in several ways. For example, if we assume $v_{it}$ and $u_{jt}$ are drawn i.i.d. from two distributions with the same mean, then $\frac{1}{T}\sum_{t}e^{-v_{it}+u_{jt}}\rightarrow 1$ by the continuous mapping theorem. Alternatively, if $X_{ij}$ is an accurate description of the long-run traffic propensity between $i,j$, then the intensity of the aggregate traffic Poisson variables $\sum_{t}Y_{ijt}$ should largely be captured by $X_{ij}$, i.e., independent of $\sum_{t}e^{-v_{it}+u_{jt}}$.
B.5 When IPF recovers the true network exactly
@@
Second, we prove that, if $X^{(t)}=A\bar{X}B$, then IPF will return $X^{(t)}$. First, if there is a finite solution to the matrix balancing problem, IPF will converge to a solution (Pukelsheim, 2014). We know that $A\bar{X}B$ is a solution, so IPF will return $D^{0}\bar{X}D^{1}$, where $(D^{0}\bar{X}D^{1})\cdot\mathbf{1}_{n}=p^{(t)}$ and $(D^{1}\bar{X}^{T}D^{0})\cdot\mathbf{1}_{m}=q^{(t)}$. Furthermore, biproportional scalings are unique with respect to marginals, meaning if two biproportional scalings $M^{1}$ and $M^{2}$ of $M^{0}$ have the same marginals, then $M^{1}=M^{2}$ (Pukelsheim, 2014). Thus, IPF will return $X^{(t)}$, since IPF will converge to $D^{0}\bar{X}D^{1}$, with marginals $p^{(t)}$ and $q^{(t)}$, and $X^{(t)}$ is the unique biproportional scaling of $\bar{X}$ that matches the marginals. ∎
@@
B.6 Poisson links uniquely recover IPF within a class of generalized linear models
@@ C.1.1 Concentration inequalities with Orlicz norms
There exists an absolute constant $C>0$ such that the following holds. Suppose that $X\sim\text{Poisson}(\lambda)$ for some $\lambda\geq 0$. Then
@@ C.2 Well-posedness of maximum likelihood estimation of biproportional Poisson
given data $X,p,q$ may not have a finite solution. Here we provide a bit more detail on this matter. The key to the well-posedness problem is the gap between the strong and the weak existence conditions discussed in Qu et al. (2023). For completeness, the conditions are:
Strong Existence. For every pair of sets of indices $I\subsetneq[m]$ and $J\subsetneq[n]$ such that $X_{ij}=0$ for $i\notin I$ and $j\in J$, $\sum_{i\in I}p_{i}\geq\sum_{j\in J}q_{j}$, with equality iff $X_{ij}=0$ for all $i\in I$ and $j\notin J$ as well.
Weak Existence. For every pair of sets of indices $I\subsetneq[n]$ and $J\subsetneq[m]$ such that $X_{ij}=0$ for $i\notin I$ and $j\in J$, $\sum_{i\in I}p_{i}\geq\sum_{j\in J}q_{j}$.
@@
We will show that the strong existence condition holds with high probability. Recall that the negative log-likelihood of the biproportional Poisson model is equivalent to
@@
Note that each $Y_{ij}$ is an independent Poisson random variable with mean and variance parameter $e^{u_{i}}X_{ij}e^{-v_{j}}$. Now for each non-empty $I\subsetneq[m]$ and $J\subsetneq[n]$ with $X_{I^{C}J}=0$, i.e., $X_{ij}\equiv 0$ for any $(i,j)$ satisfying $i\in I^{C},j\in J$, let $E_{IJ}$ be the event that $\sum_{j\in J}q_{j}=\sum_{i\in I}p_{i}$. The event $E_{IJ}$ corresponds to the case where the realized network of Poisson variables being disconnected, hence the MLE being unbounded. Our goal is to bound the union of events $E_{IJ}$ for all $I,J$ where $X_{I^{C}J}=0$. We have
@@
where in the last step we used Bennett’s inequality (Bennett, 1962). [^5] See also Terry Tao’s blog post on improved Bennett’s inequality for the Poisson distribution: https://terrytao.wordpress.com/2022/12/13/an-improvement-to-bennetts-inequality-for-the-poisson-distribution/.
@@
where we have used the fact that $X_{ij}\equiv 0$ for any $(i,j)$ satisfying $i\in I^{C},j\in J$. Now using once again that $\mathbf{1}_{m+n}$ spans the null space of $\mathcal{L}(u,v)$ for any $(u,v)$, we have
$\sum_{i\in I,j\in J^{C}}e^{u^{\ast}_{i}}X_{ij}e^{-v^{\ast}_{j}}$ $=\begin{bmatrix}e_{I}\\ e_{J}\end{bmatrix}^{T}\mathcal{L}(u^{\ast},v^{\ast})\begin{bmatrix}e_{I}\\ e_{J}\end{bmatrix}$
@@
Now applying a union bound on all $I,J$ with $X_{I^{C}J}=0$ and following the arguments from Lemma 1 in Simons & Yao (1999), we have
(MLE does not $\mathbb{P}(\text{MLE does not exist})$ $\leq\sum_{I\subsetneq[m],J\subsetneq[n],X_{I^{C}J}=0}\exp(-\frac{1}{2}\lambda_{-2}(\mathcal{L}(u^{\ast},v^{\ast}))\frac{(|I|+|J|)(m+n-|I|-|J|)}{m+n})$
@@
(MLE does not $\mathbb{P}(\text{MLE does not exist})$ $\leq 2\left[\left(1+\exp(-2\log(m+n))\right)^{m+n}-1\right]$
$=2\left[\left(1+(m+n)^{-2}\right)^{m+n}-1\right]$
@@ E.1.2 Evaluating IPF over varying sparsity in $\bar{X}$
  • The bound on the MLE’s expected estimation error, $\mathbb{E}[||(\hat{u}-u,\hat{v}-v)||^{2}_{2}]$, from Theorem 4.1, without constants:
@@
In Figure 2, we visualize our results, showing the mean and 95% CIs (from 2.5th to 97.5th percentiles over random trials) per metric over sparsity rates. We find that IPF is negatively affected by sparsity in multiple ways: the number of iterations until convergence increases, and both the expected bound and observed estimation error on the network parameters worsens. However, the rate at which these metrics change differ, with the bound growing most quickly at high levels of sparsity. To better understand why, below we analyze the connection between the bound on the MLE’s expected error, which we derived in Theorem 4.1, and the observed $\ell_{2}$ error that we evaluate in these experiments.
@@
and the right hand side qualitatively matches the shape observed in Figure 2 (right).
Sketch of concentration argument. We only include a sketch of this argument since it is standard in the random graph/matrix literature and not part of any of the main results of this work. By standard concentration estimates (i.e. Bernstein’s inequality), the number of entries of $X$ that are deleted from each row will be $rn\pm O(\sqrt{rn\log(n)})$. Equivalently, the number of entries that remain in each row will be $(1-r)n\pm O(\sqrt{rn\log(n)})$. Next we can compute that
@@ E.1.3 Testing IPF under model misspecification
Although our statistical theory is developed for MLE under correct specification, we remark that it applies more generally to non-Poisson distributions, as long as the conditional mean of $X_{ij}$ is given by $e^{u_{i}}\bar{X}_{ij}e^{-v_{j}}$. This is because pseudo-Poisson maximum likelihood is known to yield a consistent estimator as long as the logarithm of the conditional mean of the outcome is linear in the covariates. In our setting, this requirement is satisfied as long as the mean of $X_{ij}$ is of the exponential form specified in (4), without requiring the distribution to be Poisson. Since we show in Theorem 3.1 that the problem is equivalent to a Poisson regression, we can conclude that IPF yields a consistent estimate for any model with the specified log-linear means.
@@
As we did in Section E.1.2, we also test IPF over varying levels of sparsity $r$ in $\bar{X}$, but this time with $X^{(t)}$ generated from exponential distribution (38). This experiment allows us to compare the shape of $\ell_{2}$ error over varying levels of sparsity, when the data is correctly specified (generated from Poisson) vs. incorrectly specified (generated from exponential). [^9] We are not able to test the negative binomial model with sparsity greater than 0, since the number of successes $s$ must be strictly positive, so the model (40) ill-defined for $\bar{X}_{ij}=0$. We visualize the results in Figure E.3: we find that the shape of the $\ell_{2}$ curves are highly similar when the data is drawn from either distribution, and matches the expected shape of $1/\sqrt{1-r}$ (see derivation in Section E.1.2), showing the usefulness of our derived error bounds even under model misspecification.
Figure E.3: Comparing sparsity rate in $\bar{X}$ to observed $\ell_{2}$ error of IPF estimates, for data drawn from Poisson (left) vs. exponential (right) distributions. The subfigure on the left, where the model is correctly specified, exactly matches Figure 2, right. Lines represent mean and shaded region represents 95% CIs over 1000 trials.
@@ E.2.2 Convergence experiments on mobility data
Replace all zeros ConvIPF, min # edges, largest $p_{i}$ ConvIPF, min # edges, smallest $p_{i}$ ConvIPF, min change in $\lambda_{1}$
@@
# new edges $10012193$ $\mathbf{2}$ $16$ $5$
Change in $\lambda_{1}$ $30.04$ $1.85\cdot 10^{-8}$ $2.54\cdot 10^{-7}$ $\mathbf{5.60\cdot 10^{-9}}$
# ConvIPF iters. $\mathbf{2}$ $16$ $5$
@@
# new edges $10012193$ $\mathbf{2}$ $3$ $3$
Change in $\lambda_{1}$ $30.04$ $2.14\cdot 10^{-7}$ $1.02\cdot 10^{-8}$ $\mathbf{9.98\cdot 10^{-9}}$
# ConvIPF iters. $\mathbf{1}$ $2$ $2$
@@
# new edges $10012193$ $\mathbf{1}$ $9$ $5$
Change in $\lambda_{1}$ $30.04$ $1.63\cdot 10^{-8}$ $8.52\cdot 10^{-8}$ $\mathbf{5.72\cdot 10^{-9}}$
# ConvIPF iters. $\mathbf{1}$ $9$ $5$
@@ E.3 Experiments with bikeshare data
Figure E.12: Distribution of bike trips over hours.
Figure E.13: Bike trips from NYC Citibike data, from September 1 12AM to September 30 11PM, 2023.
The distribution of trips over stations is highly skewed (Figure E.13). To aggregate the individual trips into hourly network, we need to assign each trip to an hour. In 79% of trips, the trip starts and ends in the same hour, so the assignment is unambiguous. For the remaining 21% of trips, we assign the trip to the hour of its midway point. So, in our hourly network $X^{(t)}$, each $X^{(t)}_{ij}$ represents the total number of trips from station $i$ to station $j$ with its midpoint falling in hour $t$. In Figure E.13, we plot the distribution of trips over hours. We can see clear temporal patterns, e.g., fewer trips at night, two peaks on weekdays for getting to and getting from work. As usual, $p^{(t)}$ and $q^{(t)}$ are the row sums and column sums, respectively, of $X^{(t)}$. For the time-aggregated network $\bar{X}$, we explore different temporal aggregations (month, week, and day), so that we can experiment with IPF performance as the input network becomes increasingly time-aggregated.
@@ E.3.2 Convergence behavior
We observe a tradeoff between accuracy and convergence time. Using the day-aggregated network for IPF consistently outperforms the week-aggregated or month-aggregated network at recovering the true hourly network (Table E.3), but running IPF on the day-aggregated network takes substantially more iterations than using the week-aggregated network or month-aggregated network. Over the 24 hours we tested, the mean and median number of iterations when using the day-aggregated network were 7643.29 and 10000, respectively, while they were 920.79 and 419 for the week-aggregated network and 537.83 and 372 for the month-aggregated network. [^11] These are actually underestimates, since we used a maximum of 10,000 iterations. Using the month-aggregated network never reached this maximum, week-aggregated reached it in one hour out of 24, and day-aggregated reached it in most hours. This is not surprising, since the day-aggregated network is far sparser (2.04% nonzero) than the week-aggregated network (7.05% nonzero) or month-aggregated network (13.72% nonzero), and we know that IPF iterations increase with sparsity, which we also saw with synthetic data (Figure 2, left).